OpenStudy (vshiroky):

Tina Thompson scored 34 points in a recent basketball game without making any 3- point shots. She scored 23 times, making several free throws worth 1 point each and several field goals worth two points each. How many free throws did she make? How many 2 point field goals did she make?

OpenStudy (vshiroky):

Explain in terms that someone who does not know what they are doing can figure this out. Step by step.

OpenStudy (anonymous):

X+Y=23 X=23-Y REPLACE X BY (23-Y) IN THE NEXT EQUATION & SOLVE FOR Y. 1X+2Y=36 (23-Y)+2Y=36 23+Y=36 Y=36-23 Y=13 2 POINT SHOTS WERE MADE. X+13=23 X=23-13 X=10 FREE THROWS WERE MADE. PROOF: 1*10+2*13=36 10+26=36 36=36

OpenStudy (campbell_st):

well its simultaneous equations let x = field goals....and y = free throws shots x + y = 23 scoring 2x + 1y = 34 so the easiest is to use substitution... take the 1st equation and make y the subject y = 23 - x and substitute it into the scoring 2x +(23 - x) = 34 now solve for x this will be the number of 2 point shots... substitute the value into the 1st equation to find the number of free throws.

OpenStudy (vshiroky):

I don't understand how that means substitute into the scoring. .where did y go?

OpenStudy (campbell_st):

can you see how I took the shoting equation x + y = 23 and made y the subject y = 23 - x...?

OpenStudy (vshiroky):

yes

OpenStudy (vshiroky):

What happens after 2x+(23-x)=34?

OpenStudy (campbell_st):

well it becomes 2x + 23 - x = 34 does that make sense..?

OpenStudy (vshiroky):

right so x+23=34 subtract 23 from both sides making it 11 right?

OpenStudy (vshiroky):

so 11 field goals

OpenStudy (campbell_st):

thats correct... 11 from the field = 22 points... so if there were 23 shots ... how many free throws...?

OpenStudy (vshiroky):

12

OpenStudy (campbell_st):

thats it... well done